OpenStudy (ashleyy_smith):

Need help with 2-3 problems (: I'll give a medal A certain drug is made from only two ingredients: compound A and compound B. There are 7 milliliters of compound A used for every 4 milliliters of compound B. If a chemist wants to make 1089 milliliters of the drug, how many milliliters of compound B are needed?

OpenStudy (anonymous):

divide 1089 by 11

OpenStudy (anonymous):

what do you get?

OpenStudy (ashleyy_smith):

99

OpenStudy (anonymous):

ok do if you divide 1089 in to 11 parts, each part is 99 ml multiply that by 7 to get the amount of compound A, multiply 99 by 4 to get the amount of compound B

OpenStudy (ashleyy_smith):

A = 693 B = 396

OpenStudy (anonymous):

i didn't do it, would you like me to check?

OpenStudy (ashleyy_smith):

If u can (:

OpenStudy (anonymous):

i will check for you

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

btw i hope it is clear that i got the 11 from (7+4\)

OpenStudy (anonymous):

yep lol

OpenStudy (anonymous):

7+4=11 i meant

OpenStudy (anonymous):

Ashleyy_Smith will you fan me lol

OpenStudy (ashleyy_smith):

Oh okay! what about this one. Each marble bag sold by Rita's Marble Company contains 5 purple marbles for every 8 green marbles. If a bag has 45 purple marbles, how many green marbles does it contain?

OpenStudy (anonymous):

how many times does 5 go in to 45?

OpenStudy (ashleyy_smith):

9

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so each of the 9 groups of 5 purple marbles gives 8 green ones compute \(9\times 8\)

OpenStudy (ashleyy_smith):

um 72

OpenStudy (anonymous):

um right

OpenStudy (ashleyy_smith):

so that's the answer :o?

OpenStudy (ashleyy_smith):

@satellite73

OpenStudy (anonymous):

yes that is the "final answer"

OpenStudy (anonymous):

if the bag has 45 purple marbles it must have 72 green ones

OpenStudy (ashleyy_smith):

Oh yay! 2 more!

OpenStudy (ashleyy_smith):

ok it's Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 3 liters of natural oil for every 4 liters of synthetic oil. In order to make 637 liters of Petrolyn oil, how many liters of natural oil are needed?

OpenStudy (anonymous):

this is like the first one what is \(3+4\)?

OpenStudy (ashleyy_smith):

7

OpenStudy (anonymous):

ok how many times does 7 go in to 637?

OpenStudy (ashleyy_smith):

4,459

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

not 7 times 637, \(367\div7\)

OpenStudy (anonymous):

1045, is your anser

OpenStudy (ashleyy_smith):

OHH

OpenStudy (ashleyy_smith):

huh :s

OpenStudy (anonymous):

i made a typo divide \[637\div 7\]

OpenStudy (ashleyy_smith):

91

OpenStudy (anonymous):

ok so \(3\times 91\) for natural and \(4\times 91\) for synthetic

OpenStudy (anonymous):

i looked it up and i made sher is was right

OpenStudy (ashleyy_smith):

273 Natural and 364 Synthetic

OpenStudy (anonymous):

looks good

OpenStudy (ashleyy_smith):

273 right?

OpenStudy (anonymous):

yes, both are right

OpenStudy (ashleyy_smith):

I hope this is the last one.. The ratio of men to women working for a company is 5 to 6. If there are 174 women working for the company, what is the total number of employees?

OpenStudy (anonymous):

how many times does 6 go in to 174?

OpenStudy (ashleyy_smith):

29

OpenStudy (anonymous):

ok multiply that by 5 to get the number of men than add that to 174

OpenStudy (ashleyy_smith):

319 is the answer?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (ashleyy_smith):

Ok this one is the last one for sure sure

OpenStudy (ashleyy_smith):

The residents of a city voted on whether to raise property taxes. The ratio of yes votes to no votes was 5 to 3. If there were 4472 total votes, how many yes votes were there?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this is again like the first one \(5+3=8\) divide \(4472\div8\)

OpenStudy (ashleyy_smith):

559

OpenStudy (anonymous):

ok multiply that by 5 to get the number of yes votes

OpenStudy (ashleyy_smith):

2,795 woah lol

OpenStudy (anonymous):

looks good`

OpenStudy (ashleyy_smith):

Got it thank you :))

OpenStudy (anonymous):

yw, done?

OpenStudy (ashleyy_smith):

yeahhh with that module. Now I just have the other module to do :P

OpenStudy (anonymous):

have fun!!

OpenStudy (ashleyy_smith):

thanks!